OpenStudy (anonymous):

reduce to simplest terms: log2^2+log5^2

OpenStudy (anonymous):

log 100

OpenStudy (anonymous):

if your base is 10, then the answer is 2

myininaya (myininaya):

\[\log2^2+\log5^2=2\log2+2\log5=2*(\log2+\log5)\] \[=2(\log(2*5))=2(\log10)=2(1)=2 \] \[(\log_{10}(10)=1 since 10^1=10)\]